solve by elimination
y=2x+8
y=10x

Answer by waynest(281)

y = 2x + 8; y = 10x
substitute 10x for y in the first equation
10x = 2x + 8
10x - 2x = 2x - 2x + 8
8x = 8
8x/8 = 8/8
x = 1
solve for y in first and equation
y = 2(1) + 8
y = 10
check:
10 = 2(1) + 8
10 = 10
10 = 10(1)
10 = 10